OpenStudy (jewelslinda):

Object A attracts object B with a gravitational force of 5 newtons from a given distance. If the distance between the two objects is reduced in half, what will be the changed force of attraction between them?

OpenStudy (irishboy123):

\(F = \dfrac{Gm_1m_2}{d^2} = \dfrac{const }{d^2}\) \(F d^2 = const\) \(5 d^2 = F' (\dfrac{d}{2})^2\) \(F' = 20N\)